# 2022-deformable-elastic-objects-UAVs

### v1

@@ III Non-stop Flight Analysis
Now let’s consider the two orthogonal components of $\dot{\bm{f}}_{i}$ defined in (13). These vectors, together with $\dot{\bm{f}}_{i}$ and $\bm{f}_{i}$ belong all to a plane that is represented in Figure 3. In that figure, the bounds of the angles between $\dot{\bm{f}}_{i}$ and the direction of the force $\bm{f}_{i}$, given by $\bm{q}_{i}$, are highlighted by incident lines properly sloped. Moreover, the lower bound of the norm of the force derivative, $s>0$, is represented by a dotted circle; the endpoint of $\dot{\bm{f}}_{i}$ is constrained to be outside of that circle. Hence, its component on the direction orthogonal to the force, namely the one parallel to $\dot{\bm{q}}_{i}$, must lie in the green regions in Figure 3. This gives a geometrical proof that $\|\dot{\bm{q}}_{i}\|T_{i}\geq s^{\prime\prime}>0$ and so $\|\dot{\bm{q}}_{i}\|\geq\frac{s^{\prime\prime}}{T_{i}}:=s^{\prime}>0.$
@@
$\lambda_{i}=\lambda_{i}(0)+A\cos{(\psi t+\Phi_{i})}$ (19)
with constant amplitude $0<A<\infty$, initial condition $\underline{\lambda}\leq\lambda_{i}(0)\leq\overline{\lambda}$, constant phases $\Phi_{i}$ different from each other such that $|\Phi_{i}-\Phi_{j}|\geq\phi>0$ and such that $0<\underline{\Phi_{i}}\leq\Phi_{i}\leq\bar{\Phi}<\pi,$ and same positive $\psi$.
@@
Let us consider without loss of generality the force applied by cable 1. From equation (12),
@@
When ${\lambda}_{1}$ and ${\lambda}_{3}$ are chosen as in (19), they are bounded because the initial conditions and amplitudes $A$ are bounded. Especially, $\forall t,$ $|\lambda_{i}(t)|\leq|\lambda_{i}(0)|+|A|$. Moreover, their derivatives are $\dot{\lambda}_{i}=-A\psi\sin{(\psi t+\Phi_{i})}$. Because we assumed $\bm{b}_{i,j}$ are not parallel to each other, for $\|\dot{\bm{f}}_{i}\|$ to be lower bounded it is sufficient that at each time at least one of its components along the non-parallel vectors $\bm{b}_{i,j}$ is lower bounded. For the hypothesis on $\Phi_{1}$ and $\Phi_{2}$, the two terms in the RHS of (20) are such that $\forall t,$ $|\dot{\lambda}_{1}|\geq\bar{\dot{\lambda}}_{1}>0$ or ${|\dot{\lambda}|_{2}\geq\bar{\dot{\lambda}}_{2}>0}$. In the worst-case scenarios, one of the components in the RHS of (20) is zero, and $\|\dot{\bm{f}}_{i}\|=\bar{\dot{\lambda}}_{i}>0.$ Using Proposition 2, we conclude that the intensities of the carriers’ velocities are lower bounded, too. ∎
@@
(a) Three flying carriers execute non-stop flights while the suspended object pose is kept at a constant pose.
(b) One of the carriers stops during its motion because two components $\dot{\lambda}_{i}$ that generate its cable force are equal to each other in time.
(c) One of the carriers stops during its motion when the two components $\dot{\lambda}_{i}$ that generate its cable force vanish simultaneously.
Figure 5: Three different simulations. On the leftmost column, non-stop flights are shown; in the other two columns, cases in which the carriers stop. On the first row, screenshots from tasks in which the 3 flying carriers maintain the suspended object in a constant pose. The area of the object among the cable attachment points is highlighted with dotted lines, the cables are black lines, and the carriers at their endings are black dots; the path of each carrier is the colored line in blue for carrier 1, red for carrier 2, and yellow for carrier 3; the cross is the object CoM. The values of $\|\dot{\bm{p}}_{Ri}\|$ (mid row) and $\dot{{\lambda}_{i}}$ (bottom row) are reported in the corresponding columns.
@@
If two components of $\dot{\bm{\lambda}}$, let us consider without loss of generality $\dot{\lambda}_{1}$ and $\dot{\lambda}_{2}$, are such that $\dot{\lambda}_{2}=a\dot{\lambda}_{1}$, $a\in\mathbb{R}$, then the corresponding force variation, in this case as in (20), becomes
$\dot{\bm{f}}_{1}=(a+1)\dot{\lambda}_{1}(\bm{b}_{1,2}+\bm{b}_{1,3}),$ (21)
which implies that $\dot{\bm{f}}_{1}$ lies along a constant direction over time, similar to what happens for the two-carrier case. We recall that $\bm{b}_{i,j}$ are unit vectors and are not parallel to each other by hypothesis. Hence, $|\bm{b}_{1,2}+\bm{b}_{1,3}|>0$ and this case can be conducted to the two-carrier case. So, for the same reasoning of the proof of Fact 2, the velocity of carrier 1 is not lower bounded for all time, and so (10) does not hold.
@@
In any case in which $\exists t:\dot{\lambda}_{i}(t)=\dot{\lambda}_{j}(t)=0$ for some $i\neq j\in\{1,2,3\}$ then the corresponding force is zero at that time instant. For the same reasoning done in the proof of Fact 1, the corresponding carrier stops. Differently from the situation described above, the cable force variation generally is not on a straight line.
@@ IV Numerical Results
In this section, we propose simulation results that support the previously presented results.
The simulations have been carried out in Matlab-Simulink. The dynamics of each carrier is a double integrator controlled through a PD feedback law to follow a desired trajectory. Indeed, the scope of this work is to assess the feasibility of non-stop flights for the considered application. Implementation of such flights considering more realistic non-stop UAV dynamics is left for future work.
@@
Figure 5b and 5c show an example of what was referred to as degenerate case 1 and 2, respectively, in which 3 carriers maintain the object’s pose but one of them stops a few times along the path. In the simulation of case 1, $\forall i,\ \lambda_{i}(0)=2$ N $A=1$ N $\psi=1$ rad/s, $\Phi_{1}=\Phi_{3}=0$ and $\Phi_{2}=0.7$ rad. In the simulation of case 2, instead, $\forall i,\ \lambda_{i}(0)=2$ N, $A=1.2$ N, $\psi_{1}=\psi_{3}=1$ rad/s, $\psi_{2}=0.5$ rad/s, $\Phi_{1}=\Phi_{2}=0$, and $\Phi_{3}=0.7$ rad. The carriers’ velocities are reported in Figures 5e and 5f for the two cases, respectively. As it can be seen in Figures 5h and 5i, for some time instants $\dot{\lambda}_{1}=\dot{\lambda}_{j}=0$. At the same time instants, the velocity of one carrier goes to zero, too, as predicted by the theory.
@@ V Conclusions and Future Work
This work studied the compatibility of non-stop carrier flights while keeping constant the pose of cable-suspended objects. After showing that non-stop flights are not feasible when only one or two flying carriers are used, we showed that three carriers can follow non-stop flights while maintaining the object’s pose constant. We also highlighted degenerate cases in which non-stop flight is not achieved even for three carriers. Eventually, numerical examples are shown to support the theoretical analysis.
In the future, we will extend the study to a generic number of carriers $n>0$. Also, we will design an optimal planning algorithm that generates smooth non-stop trajectories for UAVs, explicitly taking into account the system’s constraints, such as the minimum forward speed of the robots, their maximum angular velocity, collision avoidance, etc. Keeping the forward speeds of the robots as constant as possible could also be considered in view of energy efficiency. More realistic simulations will be developed to assess the practical applicability of the manipulation method, e.g., considering hybrid VTOL vehicles.

### v3

@@ III Non-stop Flight Analysis
Now let’s consider the two orthogonal components of $\dot{\bm{f}}_{i}$ defined in (13), i.e. $\dot{T}_{i}\bm{q}_{i}$ and $T_{i}\dot{\bm{q}}_{i}$. These three vectors define a plane that contains also $\bm{f}_{i}=T_{i}\bm{q}_{i}$. We represent such a plane in Figure 3, where the bounds of the angles between $\dot{\bm{f}}_{i}$ and the direction of the force $\bm{f}_{i}$, given by $\bm{q}_{i}$, are represented by the dashed lines. Moreover, the lower bound of the norm of the force derivative, $s>0$, is represented by a dotted circle; the endpoint of $\dot{\bm{f}}_{i}$ is constrained to be outside of that circle. Hence, its component on the direction orthogonal to the force, namely the one parallel to $\dot{\bm{q}}_{i}$, must lie in the green regions in Figure 3, delimited by the lines at distance $\underline{v}^{\prime\prime}>0$ from the line where $\bm{f}_{i}$ lies. This gives a geometrical proof that $\|\dot{\bm{q}}_{i}\|T_{i}\geq\underline{v}^{\prime\prime}$ and so $\|\dot{\bm{q}}_{i}\|\geq\frac{\underline{v}^{\prime\prime}}{T_{i}}:=\underline{v}^{\prime}>0.$
@@
$\lambda_{i}(t)=\lambda_{i}(0)+A\cos{(\psi t+\Phi_{i})}$ (19)
where $\psi>0$; the amplitude $A$ is constant and bounded ($0<A<\infty$); the initial condition is bounded $\underline{\lambda}\leq\lambda_{i}(0)\leq\overline{\lambda}$; and the constant phases $\Phi_{i}$ are such that $0<\underline{\Phi_{i}}\leq\Phi_{i}\leq\bar{\Phi}<\pi$, and that, for every $j\neq i$ $|\Phi_{i}-\Phi_{j}|\geq\phi>0$ (non-coinciding phases).
@@
We are going to show that, choosing the aforementioned trajectories, $\|\dot{\bm{f}}_{i}\|$ is lower bounded, so we can apply Proposition 2. Let us consider without loss of generality the force applied by cable 1. From equation (12),
@@
When ${\lambda}_{1}$ and ${\lambda}_{3}$ are chosen as in (19), they are bounded because the initial conditions and amplitudes $A$ are bounded. Especially, $\forall t,$ $|\lambda_{i}(t)|\leq|\lambda_{i}(0)|+|A|$. Moreover, their derivatives are $\dot{\lambda}_{i}=-A\psi\sin{(\psi t+\Phi_{i})}$. Since $\bm{b}_{1,2}$ and $\bm{b}_{1,3}$ are not parallel, for $\|\dot{\bm{f}}_{1}\|$ to be lower bounded it is sufficient to show that at least one among $|\dot{\lambda}_{1}(t)|$ and $|\dot{\lambda}_{3}(t)|$ is non-zero at any time $t$, which is ensured by the non-coinciding phase assumption on $\Phi_{1}$ and $\Phi_{2}$. ∎
@@
(a) Three flying carriers execute non-stop flights not perturbing the pose of the suspended objetc.
(b) One of the carriers stops when two components $\dot{\lambda}_{i}$ are equal to each other.
(c) One of the carriers stopswhen the two components $\dot{\lambda}_{i}$ vanish simultaneously.
Figure 5: Results of three different simulations, one in each column. On the leftmost column, non-stop flights are shown; in the other two columns, cases in which the carriers stop. On the first row of each simulation, a frame in which the 3 flying carriers (black dots) maintain the suspended object at a constant pose. Dotted lines connect the cables’ attachment points on the object; the cables are black lines; the path of each carrier is a colored line (blue for carrier 1, red for carrier 2, and yellow for carrier 3); the cross is the object CoM. The values of $\|\dot{\bm{p}}_{Ri}\|$ and $\dot{{\lambda}_{i}}$ are also reported for each simulation.
III-D Degenerate situations in the 3-carrier case
@@
If two components of $\dot{\bm{\lambda}}$, let us consider without loss of generality $\dot{\lambda}_{1}$ and $\dot{\lambda}_{3}$, are such that $\dot{\lambda}_{3}=a\dot{\lambda}_{1}$, $a\in\mathbb{R}$, then the corresponding force variation, in this case as in (20), becomes
$\dot{\bm{f}}_{1}(t)=(a+1)\dot{\lambda}_{1}(t)(\bm{b}_{1,2}+\bm{b}_{1,3}),$ (21)
which implies that $\dot{\bm{f}}_{1}$ lies along a constant direction over time, similar to what happens for the two-carrier case. In fact, since $\bm{b}_{1,2}$ and $\bm{b}_{1,3}$ are constant vectors this case resembles the two-carrier case displayed in (17). So, for the same reasoning of the proof of Fact 2, the velocity of carrier 1 is not lower bounded for all time, and so (10) does not hold.
@@
In any case in which $\exists t:\dot{\lambda}_{i}(t)=\dot{\lambda}_{j}(t)=0$ for some $i\neq j\in\{1,2,3\}$ then the corresponding force is zero at that time instant. For the same reasoning done in the proof of Fact 1, the corresponding carrier stops. Differently from the situation described in case a), the cable force variation generally is not on a straight line.
@@ IV Numerical Results
In this section, we propose simulation results that support the previously presented results. A video of the simulation results can be found at https://youtu.be/A18oq4oTQl8. The simulations have been carried out in Matlab-Simulink. The dynamics of each carrier is a double integrator controlled through a PD feedback law to follow a desired trajectory. Indeed, the scope of this work is to assess the feasibility of non-stop flights for the considered application. The implementation considering more realistic non-stop UAV dynamics is left for future work.
@@
Figure 5b and 5c show what was referred to as degenerate cases 1 and 2, respectively: 3 carriers maintain the object’s pose but one of them stops along the path. In the simulation of case 1, $\forall i,\ \lambda_{i}(0)=2$ N $A=1$ N $\psi=1$ rad/s, $\Phi_{1}=\Phi_{3}=0$ and $\Phi_{2}=0.7$ rad. In the simulation of case 2, instead, $\forall i,\ \lambda_{i}(0)=2$ N, $A=1.2$ N, $\psi_{1}=\psi_{3}=1$ rad/s, $\psi_{2}=0.5$ rad/s, $\Phi_{1}=\Phi_{2}=0$, and $\Phi_{3}=0.7$ rad. As it can be seen in Figures 5h and 5i, for some time instants $\dot{\lambda}_{1}=\dot{\lambda}_{j}=0$. At the same time instants, the velocity of one carrier (in Figure 5e and 5f, respectively) goes to zero, too, as predicted by the theory.
@@ V Conclusions and Future Work
This work studied the possibility of non-stop carrier flights while maintaining a constant pose of cable-suspended objects. After showing that non-stop flights are not feasible for only one or two flying carriers, we showed that three carriers can follow non-stop flights while maintaining the object’s pose constant. We also highlighted degenerate cases in which non-stop flight is not achieved even for three carriers. Numerical examples supported the theoretical analysis.
In the future, we will extend the study to a generic number of carriers $n>0$. We will design an optimal planning algorithm that generates smooth non-stop trajectories for UAVs, explicitly taking into account the system’s constraints, such as the minimum forward speed of the robots, their maximum angular velocity, collision avoidance, etc. More realistic simulations will be implemented to assess the applicability of the method to VTOL hybrid platforms. Possible mechatronics adaptations to realize experiments will be considered.